OpenStudy (chrisneedshelp):

Find the value of 7 + 32 (-5 + 1) ÷ 2. 25 32 -11

OpenStudy (anonymous):

−57

OpenStudy (anonymous):

7 +32( -5 +1) /2 Remember your order of operations. We do parentheses first: 7 +32( -4) /2 Now, multiplication and division together: 7 + 32*-4 /2 7 -128 /2 7 -64 Now, subtraction. -57

OpenStudy (the_fizicx99):

Order of operations, First the parenthesis, -5 + 1 = -4 7 + 32(-4)/2 Multiplication then division, 7 -128/2 -128/2 = -64 7 + (-64) = -57 PEMDAS = Parenthesis, Exponents, Multiplication, Division, Addition and Subtraction.

OpenStudy (chrisneedshelp):

oh sorry guys its not 32, its 3^2 (2 as an exponent)

OpenStudy (anonymous):

7 +3^2( -5 +1) /2 We still do parentheses first: 7 +3^2(-4) /2 Next, exponents: 7 +9(-4) /2 Now, multiplication and division together (whichever comes first, left to right): 7 -36 /2 7 -18 Finally, addition and subtraction together (whichever comes first, left to right): -11

OpenStudy (the_fizicx99):

Oh, 7 + 3^2 (-5 + 1) ÷ 2. 7 + 3^2 * -4/2 7 + 9 * -2 7 -18 = 11

OpenStudy (chrisneedshelp):

thank you guys!!

OpenStudy (the_fizicx99):

-11, 7 is smaller than 18